OpenStudy (anonymous):

i got the answer to this question wrong A scientist counts 35 bacteria present in a culture and finds that the number of bacteria triples each hour. The function y = 35 • 3x models the number of bacteria after x hours. Graph the function. (You do not need to submit the graph). Use the graph to estimate when there will be about 550 bacteria in the culture.

OpenStudy (anonymous):

http://prntscr.com/1048q3 i tried drawing the graph

OpenStudy (anonymous):

y=35*3^x?

hartnn (hartnn):

yeah, almost correct, the point you need is the intersection point.

OpenStudy (anonymous):

\[\huge y=35\times 3x\] You have that equation. y represents the number of bacteria formed. You must find x.

hartnn (hartnn):

intersection of the horizontal line and the curve. and its 35* 3^x

OpenStudy (anonymous):

if the number of bacteria triples every hour use the function y=3*(3^x)

hartnn (hartnn):

can you find that intersection with your graph utility ?

OpenStudy (anonymous):

if you can't, you can use the desmos online graphing calculator

OpenStudy (anonymous):

http://prntscr.com/10494e

OpenStudy (anonymous):

do you need me to explain why you use y=35*x^3 ?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

every hour the number of bacteria is multiplied by 3, right?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so the first hour, there will be 35 bacteria the second hour, there will be 35*3 bacteria the third hour there will be 35*3*3 bacteria the fourth hour there will be 35*3*3*3 bacteria

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

these values are at the start of the hour, so the time values are 0,1,2 and 3 you can see that the number of bacteria each time is 35*3^0, 35*3^1, 35*3^2 and 35*3^3.

OpenStudy (anonymous):

ok then

OpenStudy (anonymous):

so using the equation y=35*3^x, what do you get as the time where y=550?